# Mean-field and fluctuations for hub dynamics in heterogeneous random networks

### v1

@@ 1.3.1 Emergent hub dynamics on power-law networks
Using the same isolated dynamics (2) and coupling function (3) as before, we fix coupling strength $\alpha=0.9$ and initialize the node states $(x_{1}^{0},x_{2}^{0},\cdots,x_{N}^{0})\in\mathbb{T}^{N}$ randomly uniformly in $[0,1)$, then iterate the $G_{1}$-network dynamics (1). We discard the first $5000$ iterates as transients and collect the next $1000$ iterates. In Figure 6 , we select three nodes of different degrees $54,875$ and $979$ for the left, middle, and right panels, respectively, and plot in red the node states $z^{t}$ against its next states $z^{t+1}$.
Note that the node behaviors vary drastically according to its degree. On the left panel, the dynamics of a node of degree 54 remain contractive; on the central panel, the node of degree 875 appears to have an expanding region; and lastly, on the right panel, the massive hub of degree 979 appears to hover around a deterministic fixed point near $z=0.2$. This shows the variety of node behaviors emergent from the interactions.
To explain, we split the coupling into contributions from low degree and non-low degree neighbors
$\sum_{j=1}^{N}A_{ij}h_{\omega_{i}^{t}}(z_{i}^{t},x_{j}^{t})=\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}h_{\omega_{i}^{t}}(z_{i}^{t},x_{j}^{t})+\sum_{j\in\mathcal{N}_{i}\setminus\mathcal{L}_{\delta}}h_{\omega_{i}^{t}}(z_{i}^{t},x_{j}^{t});$
@@
$\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\frac{1}{\nu_{i}k_{i}}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}h_{\omega_{i}^{t}}(z_{i}^{t},x_{j}^{t})\approx\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\int_{\mathbb{T}}h_{\omega_{i}^{t}}(z_{i}^{t},x)\mathrm{d}x$
@@
is the effective coupling strength that hub $i$ feels in the network dynamics, and the mean-field fluctuation of hub $i$ at time $t$ from initial datum $(\omega,x)$ is, in this example,
$\xi_{i}^{t}(\omega,x):=\frac{\alpha_{i}}{\nu_{i}k_{i}}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}\sin 2\pi x_{j}^{t}+\frac{\alpha}{\Delta_{0}}\sum_{j\in\mathcal{N}_{i}\setminus\mathcal{L}_{\delta}}\left[\sin 2\pi x_{j}^{t}-\sin 2\pi z_{i}^{t}-\frac{\omega_{i}^{t}}{3.6}\right].$
@@ 1.3.2 System size induced desynchronization
i.e., difference of states between the two most massively connected hubs $i_{0}$ of degree $\Delta_{0}=979$ and $i_{1}$ of degree $978$. For this we run the $G_{1}$-network dynamics (1) at $\alpha=0.9$ with random initial conditions, discard the first $1500$ iterated as transients, and plot for the next $1000$ iterates. Here, large values of $\eta^{t}$ indicate desynchronization. In fact, the similations in [Ric16] revealed that this desynchronization becomes rare for large system size $N$. In a statistical mechanics system with continuum symmetry, similar size induced desynchronization effects have been characterized in [BGP13].
@@ 1.3.3 Statement of main result on the power-law network
Theorem B: power-law network is locally star-like. Consider a large Chung-Lu network $G$ generated from the power-law expected degree sequence given in (6) with $N\gg 1$ nodes, power-law exponent $\beta>2$, maximum expected degree $m\asymp N^{\frac{1}{\beta-1}}$, and mean expected degree $w\ll m$. By setting hub and low degree scales
$\Delta=c_{\mathrm{hub}}m,~{}~{}c_{\mathrm{hub}}\in(0,1);~{}~{}~{}~{}\delta\asymp\Delta^{\lambda_{\mathrm{ldn}}},~{}~{}\lambda_{\mathrm{ldn}}\in(0,1),$
we regard nodes of degree above $\Delta$ as hubs and below $\delta$ as low degree. Then, with overwhelming probability we have
@@
Theorem C: hub dynamics in power-law network random dynamical system. Consider dynamics (1) on a large power-law network $G$ with $N\gg 1$ as in Theorem B. Then, with initial conditions following the uniform distribution on $\mathbb{T}^{N}$, at coupling strength $\alpha\ll\min\{N^{\lambda_{\mathrm{ldn}}\frac{\beta-2}{\beta-1}},N^{(1-\lambda_{\mathrm{ldn}})/(2\beta-2)}\}$, each hub $i\in\mathcal{H}_{\Delta}$ admits the mean-field dimensional reduction $f_{\alpha_{i},\omega}$ at effective coupling strength $\alpha_{i}$ defined to be
$f_{\alpha_{i},\omega}(z):=f_{\omega}(z)+\alpha_{i}\int_{\mathbb{T}}h(\omega,z,x)\mathrm{d}x\mod 1,~{}~{}~{}~{}\alpha_{i}:=\alpha\nu_{i}k_{i}\Delta_{0}^{-1}.$
@@ 1. $\mathrm{(i)}$
    $\mathrm{Prob}\left\{(\omega,x):\liminf_{T\to+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbm{1}_{\max\{|\xi_{i}^{t}(\omega,x)|:i\in\mathcal{H}_{\Delta}\}\leq\varepsilon}\geq 1-\exp\left(-\nu\Delta\varepsilon^{2}\alpha^{-2}/19\right)\right\}=1;$
@@ 2. $\mathrm{(ii)}$
    $\mathrm{Prob}\left\{(\omega,x):\max_{i\in\mathcal{H}_{\Delta}}\max_{t\in I_{t_{0}}^{T}}|\xi_{i}^{t}(\omega,x)|\leq 3\alpha\Delta^{-\kappa}\right\}\geq 1-\exp(-\Delta^{1-2\kappa}),~{}~{}~{}~{}0<\kappa<\min\left\{1/2,1-\lambda_{\mathrm{ldn}},\lambda_{\mathrm{ldn}}\frac{\beta-2}{\beta-1}\right\};$
@@ 3. $\mathrm{(iii)}$
    $\mathrm{Prob}\{(\omega,x):\xi_{i}^{t}(\omega,x)\leq s\}\in[F_{i}(s-s_{1})-s_{2},F_{i}(s+s_{1})+s_{2}],~{}~{}~{}~{}i\in\mathcal{H}_{\Delta},~{}~{}s\in\mathbb{R},$
@@
Main ideas of proof for Theorem C. We combine Theorems A and B. The locally star-like property from Theorem B ensures that the hub and low degree scales are separated, allowing the low degree orbit to be shadowed by an isolated orbit; again by Theorem B, most hub neighbors are of low degree and hence well-controlled by the shadowing technique from Theorem A. When excluding the bad sets for all hubs, we use Theorem B to ensure that there are only $M\sim w^{\beta-1}$ hubs.
@@ 2 Dimensional reduction
where the each node $(\Phi_{\alpha}(t,\omega,x))_{i}=x_{i}^{t}$ evolves by
$x_{i}^{t+1}=\varphi_{i}(1,\theta^{t}\omega,x_{i}^{t})+\alpha\cdot\frac{1}{\Delta_{0}}\sum_{j=1}^{N}A_{ij}h(\theta^{t}\omega,x_{i}^{t},x_{j}^{t})\mod 1,~{}~{}~{}~{}i=1,\cdots,N.$ (8)
Note that the trajectory $\{x_{j}^{t}:t\in\mathbb{N}\}$ of low degree node $j\in\mathcal{L}_{\delta}$ is a pseudo-orbit of $\varphi_{j}$ with error at each time step bounded by $\alpha\Delta_{0}^{-1}\delta\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}$.
@@
    Independent and identically distributed node maps: $\omega\mapsto\varphi_{i}(1,\theta^{t}\omega,\cdot)$, $i=1,\cdots,N$, $t\in\mathbb{N}$ share the same distribution in the space of continuous circle maps and are independent in time $t$ and node $i$;
@@ 2.1 Reduction on the star network
$|\xi_{m}(\theta^{t}\omega,\Phi_{\alpha}(t,\omega,x))|\leq\varepsilon.$
Given a real sequence $\{\xi(t)\}$, we say that $\xi(t)>\varepsilon$ occurs with asymptotic frequency at most $\rho$ if
$\limsup_{T\to+\infty}\frac{1}{T}\#\left\{t\in[0,T-1]:\xi(t)>\varepsilon\right\}\leq\rho.$
@@
Suppose (R1–3) hold for the star network random dynamical system (9-10) on $L=N-1$ low-degree nodes and one hub node at coupling strength $\alpha>0$. Let initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ be such that the network trajectory $\{x^{t}=\Phi_{\alpha}(t,\omega,x):t\in\mathbb{N}\}$ admits an $m$-typical shadowing orbit in the low degree coordinates; more precisely, there is $(\omega_{{s}},x_{{s}})\in\Omega\times\mathbb{T}^{N}$ satisfying
$d_{\mathbb{T}}(x_{j}^{t},\varphi_{j}(t,\omega_{{s}},x_{{s},j}))\leq\varepsilon_{s}\left(\alpha L^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right),~{}~{}~{}~{}\forall t\in\mathbb{N},\forall j=2,\cdots,N;$ (Shadowing)
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi_{j}(t,\omega_{{s}},x_{{s},j})}\xrightarrow[T\to+\infty]{*}m,~{}~{}~{}~{}\forall j=2,\cdots,N,$ (Typicality)
where $d_{\mathbb{T}}$ denotes the distance on the circle, the shadowing precision $\varepsilon_{s}(\delta_{s})$ is a $\mathbb{R}_{+}$-valued function converging to 0 as $\delta_{s}$ tends to 0. Then, for any error tolerance
@@
$\rho(\varepsilon,\omega)=D(\varepsilon,\omega)\exp(-L\varepsilon^{2}\alpha^{-2}c(\omega)),~{}~{}~{}~{}D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}},~{}~{}c(\omega)=\frac{450}{\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}},$
@@
(Typicality) condition assumes that the shadowing $\varphi_{j}$-orbit $\varphi_{j}(t,\omega_{{s}},x_{{s},j})$ is $m$-typical, i.e., follows the distribution $m$ given by the unique stationary meausre. Note that Theorem 2.3 (Shadowing) plus (Typicality) for one initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ and concludes the reduction for this particular instance of initial data. To obtain Shadowing and Typicality conditions for $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ as in Theorem A, one tends to encounter the following scenario. By local hyperbolicity of node dynamics $\varphi_{i}$, the shadowing technique produces some initial datum $(\omega_{s},x_{s})$, whose random orbit $\varphi_{i}(t,\omega_{s},x_{s})$ achieves (Shadowing). And we wish to establish (Typicality) for $\varphi_{i}(t,\omega_{s},x_{s})$. For this, the existence and uniqueness of stationary measure $m$ from (R2) ensures, see [BW22, Theorem 16.4], the ergodicity of the Markov chain associated to the iid random iteration $\varphi_{i}$ on each node $i=1,\cdots,N$ with initial distribution $m$.
@@
    By Breiman’s ergodic theorem [Bre60], for each $x_{i}\in\mathbb{T}$, there is $\Omega_{x_{i}}\subseteq\Omega$ with $\mathbb{P}(\Omega_{x_{i}})=1$ such that the random orbit $\varphi_{i}(t,\omega,x_{i})$ is $m$-typical for every $\omega\in\Omega_{x_{i}}$. However, this typicality falls short because there may be pathological situations where the shadowing initial datum $(\omega_{s},x_{s,i})$ is such that $\omega_{s}\notin\Omega_{x_{s,i}}$.
@@
    Using the uniform contraction property as in the examples from Introduction, together with Breiman’s ergodic theorem, see Theorem 3.3, we obtain a uniform $\Omega_{*}$ with $\mathbb{P}(\Omega_{*})=1$ such that the random orbit $\varphi_{i}(t,\omega,x_{i})$ is $m$-typical for every $(\omega,x_{i})\in\Omega_{*}\times\mathbb{T}$. Moreover, we show that $(\omega_{s},x_{s,i})=(\omega,x_{i})$, see Corollary 3.5. Thus, Shadowing and Typicality conditions are achieved for every $(\omega,x)\in\Omega_{*}\times\mathbb{T}^{N}$.
@@
If $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ admits shadowing intial data $(\omega_{{s}},x_{{s}})$ satisfying (Shadowing) and (Typicality), then the $\varepsilon$-reduction holds $\mathbb{P}\otimes m^{\otimes N}$-almost surely by Theorem 2.3. In this case, if there is an “absolutely continuous shadowing map" $(\omega,x)\mapsto(\omega_{{s}},x_{{s}})$, then one obtains (ii) Small fluctuation in long time windows and (iii) Gaussian fluctuations, as in Theorem A. See Appendix A for more details.

### v2

@@ 1.3.1 Emergent hub dynamics on power-law networks
Using the same isolated dynamics (2) and coupling function (3) as in Example 1.2, we fix coupling strength $\alpha=0.9$ and initialize the node states $(x_{1}^{0},x_{2}^{0},\cdots,x_{N}^{0})\in\mathbb{T}^{N}$ randomly uniformly in $[0,1)$, then iterate the $G_{1}$-network dynamics (1). We discard the first $5000$ iterates as transients and collect the next $1000$ iterates. In Figure 6 , we select three nodes of different degrees $54,875$ and $979$ for the left, middle, and right panels, respectively, and plot in red the hub states $z^{t}$ against its next states $z^{t+1}$.
Note that the node behaviors vary drastically according to their degree. On the left panel, the dynamics of a node of degree 54 remain contractive; on the central panel, the node of degree 875 appears to have an expanding region; and lastly, on the right panel, the massive hub of degree 979 appears to hover around a deterministic fixed point near $z=0.2$. This shows the variety of node behaviors emergent from the interactions.
To explain, we continue to write $z_{i}^{t}$ for $x_{i}^{t}$ to emphasize the hubs $i\in\mathcal{H}_{\Delta}$. Split the coupling into contributions from low degree and non-low degree neighbors
$\sum_{j=1}^{N}A_{ij}h_{\boldsymbol{\omega}_{i}^{t}}(z_{i}^{t},x_{j}^{t})=\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}h_{\boldsymbol{\omega}_{i}^{t}}(z_{i}^{t},x_{j}^{t})+\sum_{j\in\mathcal{N}_{i}\setminus\mathcal{L}_{\delta}}h_{\boldsymbol{\omega}_{i}^{t}}(z_{i}^{t},x_{j}^{t});$
@@
$\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\frac{1}{\nu_{i}k_{i}}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}h_{\boldsymbol{\omega}_{i}^{t}}(z_{i}^{t},x_{j}^{t})\approx\alpha\frac{\nu_{i}k_{i}}{\Delta_{0}}\int_{\mathbb{T}}h_{\boldsymbol{\omega}_{i}^{t}}(z_{i}^{t},x)\mathrm{d}x$
@@
is the effective coupling strength that hub $i$ feels in the network dynamics, and the mean-field fluctuation of hub $i$ at time $t$ from initial datum $(\boldsymbol{\omega},x)$ is, in this example,
$\xi_{i}^{t}(\boldsymbol{\omega},x):=\frac{\alpha_{i}}{\nu_{i}k_{i}}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{L}_{\delta}}\sin 2\pi x_{j}^{t}+\frac{\alpha}{\Delta_{0}}\sum_{j\in\mathcal{N}_{i}\setminus\mathcal{L}_{\delta}}\left[\sin 2\pi x_{j}^{t}-\sin 2\pi z_{i}^{t}-\frac{\boldsymbol{\omega}_{i}^{t}}{3.6}\right].$
@@ 1.3.2 System size induced desynchronization
i.e., difference of states between the two most massively connected hubs $i_{0}$ of degree $\Delta_{0}=979$ and $i_{1}$ of degree $978$. For this we run the $G_{1}$-network dynamics (1) at $\alpha=0.9$ with random initial conditions, discard the first $1500$ iterated as transients, and plot for the next $1000$ iterates. Here, large values of $\eta^{t}$ indicate desynchronization. In fact, the simulations in [Ric16] revealed that this desynchronization becomes rare for large system size $N$. In a statistical mechanics system with continuum symmetry, similar size induced desynchronization effects have been characterized in [BGP13].
@@ 1.3.3 Statement of main result on the power-law network
Theorem B: power-law network is locally star-like. Fix parameters $c_{\mathrm{hub}},\lambda_{\mathrm{ldn}}\in(0,1)$. Consider a large Chung-Lu network $G$ generated from the power-law expected degree sequence given in (6) with $N$ nodes, power-law exponent $\beta>2$, maximum expected degree $m\asymp N^{\frac{1}{\beta-1}}$, and mean expected degree $w=o(N^{\lambda_{\mathrm{ldn}}\frac{1}{\beta-1}})$. By setting hub and low degree scales
$\Delta=c_{\mathrm{hub}}m,~{}~{}~{}~{}\delta\asymp\Delta^{\lambda_{\mathrm{ldn}}},$
we regard nodes of degree above $\Delta$ as hubs and below $\delta$ as low degree. Then, with probability $1-O(N^{-1/5})$, we have
@@
Theorem C: hub dynamics in power-law network random dynamical system. Consider dynamics (1) on a large power-law network $G$ as in Theorem B and $f_{\omega},h_{\omega}$ as in Example 1.2. Then, with initial conditions following the uniform distribution on $\mathbb{T}^{N}$, at coupling strength $\alpha=o\left(\min\{N^{\lambda_{\mathrm{ldn}}\frac{\beta-2}{\beta-1}},N^{(1-\lambda_{\mathrm{ldn}})/(2\beta-2)}\}\right)$, each hub $i\in\mathcal{H}_{\Delta}$ admits the mean-field dimensional reduction $f_{\alpha_{i},\omega}$ at effective coupling strength $\alpha_{i}$ defined to be
$f_{\alpha_{i},\omega}(z):=f_{\omega}(z)+\alpha_{i}\int_{\mathbb{T}}h_{\omega}(z,x)\mathrm{d}x\mod 1,~{}~{}~{}~{}\alpha_{i}:=\alpha\nu_{i}k_{i}\Delta_{0}^{-1}.$
More precisely, we have:
@@ 1. $\mathrm{(i)}$
    $\mathrm{Prob}\left\{(\boldsymbol{\omega},x):\liminf_{T\to+\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbbm{1}_{\max\{|\xi_{i}^{t}(\boldsymbol{\omega},x)|:i\in\mathcal{H}_{\Delta}\}\leq\varepsilon}\geq 1-\exp\left(-\nu\Delta\varepsilon^{2}\alpha^{-2}/19\right)\right\}=1;$
@@ 2. $\mathrm{(ii)}$
    $\mathrm{Prob}\left\{(\boldsymbol{\omega},x):\max_{i\in\mathcal{H}_{\Delta}}\max_{t\in I_{t_{0}}^{T}}|\xi_{i}^{t}(\boldsymbol{\omega},x)|\leq 3\alpha\Delta^{-\kappa}\right\}\geq 1-\exp(-\Delta^{1-2\kappa}),~{}~{}~{}~{}0<\kappa<\min\left\{1/2,1-\lambda_{\mathrm{ldn}},\lambda_{\mathrm{ldn}}\frac{\beta-2}{\beta-1}\right\};$
@@ 3. $\mathrm{(iii)}$
    $\mathrm{Prob}\{(\boldsymbol{\omega},x):\xi_{i}^{t}(\boldsymbol{\omega},x)\leq s\}\in[F_{i}(s-s_{1})-s_{2},F_{i}(s+s_{1})+s_{2}],~{}~{}~{}~{}i\in\mathcal{H}_{\Delta},~{}~{}s\in\mathbb{R},$
@@
Main ideas of proof for Theorem C. We combine Theorems A and B. The locally star-like property from Theorem B ensures that the hub and low degree scales are separated, allowing the low degree orbit to be shadowed by an isolated orbit; again by Theorem B, most hub neighbors are of low degree and hence well-controlled by the shadowing technique from Theorem A. When excluding the bad sets for all hubs, we use Theorem B to ensure that there are only $M\sim w^{\beta-1}$ hubs. The detailed proof of Theorem C can be found in Appendix A.2.
@@ 2 Dimensional reduction
where each node $(\Phi_{\alpha}(t,\omega,x))_{i}=x_{i}^{t}$ evolves by
$x_{i}^{t+1}=\varphi_{i}(1,\theta^{t}\omega,x_{i}^{t})+\alpha\cdot\frac{1}{\Delta_{0}}\sum_{j=1}^{N}A_{ij}h(\theta^{t}\omega,x_{i}^{t},x_{j}^{t})\mod 1,~{}~{}~{}~{}i=1,\cdots,N,$ (8)
where $A=(A_{ij})_{i,j}$ is the same adjacency matrix of network $G$ as in Eq. (1). Note that the trajectory $\{x_{j}^{t}:t\in\mathbb{N}\}$ of low degree node $j\in\mathcal{L}_{\delta}$ is a pseudo-orbit of $\varphi_{j}$ with error at each time step bounded by $\alpha\Delta_{0}^{-1}\delta\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}$.
@@
    Independent and identically distributed node maps: the random variables $\omega\mapsto\varphi_{i}(1,\theta^{t}\omega,\cdot)$, $i=1,\cdots,N$, $t\in\mathbb{N}$ take values in the space of continuous circle maps, have the same distribution, and are independent in time $t$ and node $i$;
@@
Remark 2.2 (Notation of $\omega$).
To avoid notational cluter, we have changed the notation of noise. Now we only use the simple font $\omega$ as an element of the abstract probability space $\Omega$ to denote noise realization. The fact that the noise is iid in time $t=0,1,\cdots$ and node coordinates $i=1,\cdots,N$ is reflected in node dynamics $\varphi_{i}(t,\omega,\cdot)$ and coupling function $h(\theta^{t}\omega,\cdot,\cdot)$.
@@ 2.1 Reduction on the star network
$|\xi_{m}(\theta^{t}\omega,\Phi_{\alpha}(t,\omega,x))|\leq\varepsilon,$
and that it admits $\varepsilon$-reduction to $\varphi_{\alpha,m}$ on initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ with exceptional frequency at most $\rho$ if
$\limsup_{T\to+\infty}\frac{1}{T}\#\left\{t\in[0,T-1]:|\xi_{m}(\theta^{t}\omega,\Phi_{\alpha}(t,\omega,x))|>\varepsilon\right\}\leq\rho.$
@@
Suppose (R1–3) hold for the star network random dynamical system (9-10) on $N-1$ low-degree nodes and one hub node at coupling strength $\alpha>0$. Let initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ be such that the network trajectory $\{x^{t}=\Phi_{\alpha}(t,\omega,x):t\in\mathbb{N}\}$ admits an $m$-typical shadowing orbit in the low degree coordinates; more precisely, there is $(\omega_{{s}},x_{{s}})\in\Omega\times\mathbb{T}^{N}$ satisfying
$\sup_{t\in\mathbb{N}}\max_{j=2,\cdots,N}d_{\mathbb{T}}(x_{j}^{t},\varphi_{j}(t,\omega_{{s}},x_{{s},j}))\leq\varepsilon_{s}\left(\alpha(N-1)^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right);$ (Shadowing)
$\frac{1}{T}\sum_{t=0}^{T-1}\delta_{\varphi_{2}(t,\omega_{{s}},x_{{s},2})}\otimes\cdots\otimes\delta_{\varphi_{N}(t,\omega_{{s}},x_{{s},N})}\xrightarrow[T\to+\infty]{\text{weak}^{*}}m^{\otimes(N-1)},$ (Typicality)
where $d_{\mathbb{T}}$ denotes the distance on the circle, the shadowing precision $\delta_{s}\mapsto\varepsilon_{s}(\delta_{s})$ is a $\mathbb{R}_{+}$-valued function converging to 0 as $\delta_{s}$ tends to 0. Then, for any error tolerance
$\varepsilon\geq\max\left\{\alpha(N-1)^{-1/2},4\alpha\sup_{t\in\mathbb{N}}|h(\theta^{t}\omega,\cdot,\cdot)|_{\mathrm{Lip}}\varepsilon_{s}\left(\alpha(N-1)^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{0}}\right)\right\},$
@@
$\rho(\varepsilon,\omega)=D(\varepsilon,\omega)\exp(-(N-1)\varepsilon^{2}\alpha^{-2}c(\omega)),~{}~{}~{}~{}D(\varepsilon,\omega)\asymp\varepsilon^{-1}\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}},~{}~{}c(\omega)=\frac{450}{\sup_{t\in\mathbb{N}}\|h(\theta^{t}\omega,\cdot,\cdot)\|_{C^{4}}},$
@@
(Typicality) condition assumes that the shadowing orbit $(\varphi_{2}(t,\omega_{{s}},x_{{s},2}),\cdots,\varphi_{N}(t,\omega_{{s}},x_{{s},N}))$ is $m^{\otimes L}$-typical, $L=N-1$, i.e., follows the distribution $m^{\otimes L}$ given by the unique stationary meausre. Note that Theorem 2.4 (Shadowing) plus (Typicality) for one initial data $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ and concludes the reduction for this particular instance of initial data. To obtain Shadowing and Typicality conditions for $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ as in Theorem A, one tends to encounter the following scenario. By local hyperbolicity of node dynamics $\varphi_{i}$, the shadowing technique produces some initial datum $(\omega_{s},x_{s})$, whose random orbit $\varphi_{i}(t,\omega_{s},x_{s})$ achieves (Shadowing). And we wish to establish (Typicality) for $\varphi_{i}(t,\omega_{s},x_{s})$. For this, the existence and uniqueness of stationary measure $m$ from (R2) ensures, see [BW22, Theorem 16.4], the ergodicity of the Markov chain associated to the iid random iteration $\varphi_{i}$ on each node $i=1,\cdots,N$ with initial distribution $m$.
@@
    By Breiman’s ergodic theorem [Bre60], for each $x_{i}\in\mathbb{T}$, there is $\Omega_{x_{i}}\subseteq\Omega$ with $\mathbb{P}(\Omega_{x_{i}})=1$ such that the random orbit $\varphi_{i}(t,\omega,x_{i})$ is $m$-typical for every $\omega\in\Omega_{x_{i}}$. However, this typicality falls short because there may be pathological situations where the shadowing initial datum $(\omega_{s},x_{s,i})$ is such that $\omega_{s}\notin\Omega_{x_{s,i}}$. This is the case with random expanding or hyperbolic maps as node dynamics.
@@
    For random uniform contractions, such as those in Example 1.2, Theorem 3.3 establishes a uniform $\Omega_{*}$ with $\mathbb{P}(\Omega_{*})=1$ such that the random orbit $\varphi_{i}(t,\omega,x_{i})$ is $m$-typical for every $(\omega,x_{i})\in\Omega_{*}\times\mathbb{T}$. Moreover, we show that $(\omega_{s},x_{s})=(\omega,x)$, see Corollary 3.5. Thus, Shadowing and Typicality conditions are achieved for every $(\omega,x)\in\Omega_{*}\times\mathbb{T}^{N}$.
@@
If $\mathbb{P}\otimes m^{\otimes N}$-a.e. $(\omega,x)\in\Omega\times\mathbb{T}^{N}$ admits shadowing intial data $(\omega_{{s}},x_{{s}})$ satisfying (Shadowing) and (Typicality), then the $\varepsilon$-reduction holds $\mathbb{P}\otimes m^{\otimes N}$-almost surely by Theorem 2.4. In the case of Example 1.2, we have $(\omega,x)=(\omega_{{s}},x_{{s}})$ and hence obtain (ii) Small fluctuation in long time windows and (iii) Gaussian fluctuations, as in Theorem A. See Appendix A for more details.
Remark 2.7.
As in the usual Shadowing Lemmas, $\epsilon_{s}(\delta_{s})$ in Theorem 2.4 is the shadowing precision which upper bounds the distance between the shadowing and pseudo orbits; it is a function of the error tolerance $\delta_{s}$ of the pseudo orbit. In Example 1.2, we have $\delta_{s}=\alpha L^{-1}$ and $\epsilon_{s}(\delta_{s})=\frac{\delta_{s}}{1-\lambda}=\frac{\alpha L^{-1}}{1-\lambda}$, where $\lambda=1/2$ is the contraction rate, see also Corollary 3.5. Our result relies on the hyperbolicity of the low degree node dynamics alone in terms of the shadowing property, which allows us to overcome the shortcoming i) of [PvST20] as discussed in Introduction.